OpenStudy (anonymous):

How many bows each using two thirds of a yard of ribbon can be made from 20 yards of ribbon?

OpenStudy (anonymous):

I THINK you do set up the problem like this: \[\frac{ 2}{ 3 }\times20\]can anyone confirm?

OpenStudy (anonymous):

\[\frac{20}{\frac{2}{3}}=30 \]

OpenStudy (anonymous):

If each bow took 1 yard, then at least 20 bows could be made.